### v1

# Graph Masked Autoencoder for Spatio-Temporal Graph Learning

Abstract
Effective spatio-temporal prediction frameworks play a crucial role in urban sensing applications, including traffic analysis, human mobility behavior modeling, and citywide crime prediction. However, the presence of data noise and label sparsity in spatio-temporal data presents significant challenges for existing neural network models in learning effective and robust region representations. To address these challenges, we propose a novel spatio-temporal graph masked autoencoder paradigm that explores generative self-supervised learning for effective spatio-temporal data augmentation. Our proposed framework introduces a spatial-temporal heterogeneous graph neural encoder that captures region-wise dependencies from heterogeneous data sources, enabling the modeling of diverse spatial dependencies. In our spatio-temporal self-supervised learning paradigm, we incorporate a masked autoencoding mechanism on node representations and structures. This mechanism automatically distills heterogeneous spatio-temporal dependencies across regions over time, enhancing the learning process of dynamic region-wise spatial correlations. To validate the effectiveness of our STGMAE framework, we conduct extensive experiments on various spatio-temporal mining tasks. We compare our approach against state-of-the-art baselines. The results of these evaluations demonstrate the superiority of our proposed framework in terms of performance and its ability to address the challenges of spatial and temporal data noise and sparsity in practical urban sensing scenarios.
Index Terms:
Graph Autoencoder, Spatio-Temporal Graph Learning
I Introduction
As remote sensing technologies and large-scale computing infrastructures continue to provide an abundance of spatio-temporal data, there is an increasing demand for robust spatio-temporal prediction frameworks in urban sensing applications. These applications include traffic analysis, such as the works by STGODE,  [1, 2, 3, 4] as well as modeling human mobility behavior, as explored by studies  [5, 6, 7]. Additionally, citywide crime prediction, as studied by Chen and Diao [8, 9], is another important area of focus. A crucial aspect of developing effective spatio-temporal prediction frameworks lies in the accurate capture of spatial and temporal correlations between different geographical locations and time intervals. This ability ensures comprehensive insights into dynamic urban phenomena and supports informed decision-making processes.
Researchers have introduced spatio-temporal graph neural networks (STGNNs) as a promising learning paradigm. Notable works in this area include [10, 11, 12]. These networks are specifically designed to handle data with graph structures and have garnered significant attention due to their ability to capture multi-hop spatial and temporal dependencies. This is accomplished through the iterative propagation of messages among nodes, leveraging their knowledge of both spatial and temporal information. By employing this recursive message passing mechanism, STGNNs demonstrate exceptional proficiency in modeling intricate dependencies in spatio-temporal data. Their effectiveness lies in their capability to integrate and exploit the inherent structure and temporal dynamics of the data, enabling more accurate predictions and insightful analysis of complex spatio-temporal phenomena in various real-world applications.
However, existing works in STGNNs have not adequately addressed the challenges of noise and sparsity in spatio-temporal data, which hinders the accurate modeling of dependencies. For instance, traffic data collected from sensors or GPS devices may contain outliers or missing values, leading to less relevant region connections and noisy signals. These noisy signals can negatively impact the performance of spatio-temporal prediction tasks, as they introduce irrelevant information and hinder the accurate modeling of dependencies. Besides, most existing solutions approach spatio-temporal prediction tasks in a supervised manner, relying on a sufficient amount of labeled data for training. However, spatio-temporal data is often sparsely labeled, especially for rare urban anomalies or events. This sparsity presents a significant challenge in building accurate prediction models. Thus, it becomes crucial to develop learning models that can effectively handle sparse labeled data and learn from limited samples for achieving accurate spatio-temporal predictions in real-world urban scenarios.
In this paper, we propose Spatio-Temporal Graph Masked Autoencoder (STGMAE) to address the challenges of noise and sparsity in spatio-temporal data. Our framework leverages multi-view spatio-temporal data and introduces a spatio-temporal heterogeneous graph neural encoder to capture region-wise dependencies. By considering both intra-view and inter-view data correlations, our model effectively models spatio-temporal patterns. Moreover, we propose a self-supervised learning paradigm that utilizes the mask mechanism to distill heterogeneous dependencies across observed regions in urban space. Specifically, we adopt the masked autoencoding mechanism on the node representations and graph structures to identify and emphasize the most relevant connections and features for prediction tasks. This enhances the learning process of implicit cross-region relations and improves the robustness of the model against noisy signals.
At the core of our approach is a generative spatio-temporal graph auto-encoder. This model learns to reconstruct both node features and the graph structure by utilizing the mask mechanism. By selectively masking certain regions in the input data during training, our model focuses on capturing the dependencies between regions and effectively reconstructing the missing or corrupted features and connections. This capability effectively overcomes the data sparsity by providing accurate self-supervisory signals to prioritize the most relevant information for spatio-temporal relation learning.
Our STGMAE framework has been extensively evaluated in diverse spatio-temporal mining tasks and compared against multiple region representation and GNN baselines. The results of these evaluations unequivocally demonstrate the effectiveness and superiority of our proposed method in addressing the challenges associated with data noise. We highlight the key contributions as follows:
  • We place a strong emphasis on addressing the challenges of data noise in spatio-temporal prediction tasks. To achieve this, we introduce the superiority of generative contrastive learning in region representation, incorporating the mask mechanism to enhance the robustness and generalization ability of our spatio-temporal heterogeneous graph neural architecture.
  • We propose a spatio-temporal heterogeneous graph neural network that captures both intra-view and inter-view region dependencies while modeling various spatio-temporal factors. To address the data sparsity issue, we perform generative spatio-temporal self-supervised learning over the region dependency graph. Our framework is built upon the masked autoencoding mechanism via masking node features and the graph structure on the heterogeneous spatio-temporal graph.
  • We rigorously validate our STGMAE across a range of spatio-temporal mining tasks, conducting comprehensive comparisons with state-of-the-art baselines. Our method consistently achieves the best performance across all tasks, including crime prediction and traffic prediction.
II Related Work
Self-supervised Learning for Spatio-Temporal Graphs. In the field of self-supervised learning (SSL) for spatio-temporal graphs, research has focused on capturing spatial and temporal dependencies within the graph structure. Recent studies [13, 14, 15, 16] propose innovative approaches in this area. Liu et al. [13] address data sparsity in POI recommendation by leveraging supplementary information through SSL. Ji et al. [14] enhance traffic prediction with SSL, capturing spatial and temporal heterogeneity. Zhang et al. [15] improve region representations using SSL, capturing underlying patterns and dynamics. Zhang et al. [16] combine SSL with adversarial training, enhancing graph representation and downstream task performance. These studies demonstrate the effectiveness of SSL in capturing spatio-temporal dependencies and improving representation learning for various tasks such as POI recommendation, traffic prediction, and region representations.
Graph Mask Autoencoder. Graph autoencoder methods are generative self-supervised learning approaches that construct input graphs directly, eliminating the need for data augmentation. Kipf and Welling [17] introduce Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) using Graph Convolutional Networks (GCN) for topology and node content reconstruction. ARVGA [18] proposes an adversarial graph embedding framework, enhancing robustness and generalization. MGAE [19] and GALA [20] combine structure and node reconstruction for a comprehensive graph representation. GraphMAE [21] focuses on node reconstruction, achieving promising results in node classification. Overall, graph autoencoders offer a powerful framework for self-supervised learning on graphs, excelling in tasks like graph generation and node classification.
Region Representation Learning. Research on the region representation learning problem includes the following studies [22, 23, 24, 25, 26, 27, 15]. In particular, Fu et al. [26] suggest to incorporate both intra-region information (such as POI distance inside a region) and inter-region information (such as the similarity of POI distributions between two regions) to increase the quality of region representations. The concept of putting forward in [26] is expanded upon by Zhang et al. [25] by using a collective adversarial training.
In a recent study conducted by Zhang et al. [24], a multi-view joint learning model is developed to learn region representations. To capture region correlations from different perspectives, such as human mobility and region properties, the study incorporates a graph attention technique for each view. However, this approach focuses solely on modeling region correlations and does not consider the inclusion of point-of-interest (POI) data, which is essential for capturing region functionality. Similarly, Wu et al. [27] propose a method for learning area representations by extracting traffic patterns. However, their approach also excludes POI data and relies solely on mobility data. The efficacy of these techniques heavily relies on the generation of high-quality region graphs. Furthermore, they may struggle to learn accurate region representations when faced with noisy and skewed spatio-temporal data. In contrast, a recent study by Zhang et al. [15] proposes the adoption of a contrastive self-supervised learning paradigm on a multi-view heterogeneous spatio-temporal graph. This approach utilizes a self-contrastive learning method to obtain high-quality region vectors. By contrast, our paper employs a generative self-supervised learning method to model region correlations by masking links and nodes on the heterogeneous spatio-temporal graph.
III Methodology
Figure 1: Architecture of our STGMAE region representation model with graph masked autoencoder.
The framework of STGMAE is depicted in Figure 1, which addresses the task of reconstructing links and features of masked nodes using a mask mechanism based on generative self-supervised learning on the heterogeneous spatio-temporal graph. The STGMAE framework consists of an encoder and a decoder. The encoder utilizes GCN [28] to capture the underlying graph structure and extract meaningful representations from the heterogeneous spatio-temporal graph $\mathcal{G}$. The decoder, also employing GCN, is responsible for reconstructing the masked links and features of nodes in the graph. The STGMAE framework operates on the heterogeneous spatio-temporal graph $\mathcal{G}$, which incorporates various types of information, such as spatial, temporal, and attribute data. By utilizing the mask mechanism, STGMAE learns to reconstruct the missing links and features of masked nodes, thereby capturing the correlations and dependencies. And the specific algorithm is shown in Algorithm 1.
III-A Preliminaries
In our methodology, we employ a partitioning approach to divide the geographic area of a city (such as New York) into $I$ spatio-regions. Each spatio-region is identified by an index, denoted as $i$ (e.g., $r_{i}$). To capture a wide range of urban contextual information, we propose a region embedding framework that integrates multiple diverse data sources associated with the metropolitan space. These data sources serve as inputs to our STGMAE, and enriched representation of each spatio-region.
Region Point-of-Interests (POIs). We create a POI matrix $\mathbf{P}\in{\mathbb{R}^{I\times C}}$, where $C$ is the number of POI categories, to characterize the region’s urban functions (such as entertainment and health).
Human Mobility Trajectories. We create a matrix $\mathbf{M}\in\mathbb{R}^{I\times I}$ utilizing human trajectory to accurately represent the dynamic urban flows between citywide zones, where $T$ is the number of time slots. Individual human mobility traces are saved in this collection as $(r_{s},r_{d},t_{s},t_{d})$, where $r_{s}$ and $r_{d}$ stand in for source and destination regions, respectively, and $t_{s}$ and $t_{d}$ stand for timestamp data.
Spatio Correlations. We create a spatio matrix $\mathcal{D}\in\mathbb{R}^{I\times T}$. It is composed of region pairs like $(r_{s},r_{d},d^{\prime})$ whose geographical distances are less than a threshold $\epsilon$, where $d^{\prime}$ denotes the distance from the source node $r_{s}$ to the destination node $r_{d}$.
Problem Statement: Given the spatio-temporal graph, denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are node set and edge set respectively. $\mathcal{G}$ is built upon the POI matrix $\mathbf{P}$, the mobility matrix $\mathbf{M}$ and the spatio matrix $\mathcal{D}$. The task of this work is to learn a function $f$, which maps each node of $\mathcal{G}$ to the embedding space and reconstruct both the structure $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and features $\mathbf{X}\in\mathbb{R}^{N\times D}$ of the input spatio-temporal graph $\mathcal{G}$. Meanwhile, the embedding space can effectively preserve the geographical and temporal connections between various locations and time slots, which is advantageous for a variety of urban sensing applications (e.g.traffic forecasting and crime prediction).
III-B Embedding Layer
To incorporate the Point-of-Interest (POI) context into region embeddings, we have developed a POI context embedding layer that reflects the functional information of each region in the latent representation space. Inspired by the POI token embedding technique using Skip-gram [29], we input the region-specific POI vector into the Skip-gram model [30] for POI context embedding. Subsequently, we concatenate the region-specific POI embeddings to generate POI-aware representations: $\bar{\textbf{E}}=\text{MLP}(\text{Skip-gram}(\mathcal{P}))$. Here, we employ a Multi-Layer Perceptron (MLP) for the embedding projection. The embedding table $\textbf{E}\in\mathbb{R}^{I\times d}$ represents the embeddings for all regions, where $d$ is the hidden dimensionality. Each region’s embedding $\textbf{e}_{i}$ captures the region-specific POI contextual information in the latent semantic space. To incorporate region interactions within our POI embedding layer, we introduce a region-wise self-attention mechanism given by:
$\textbf{e}_{i}=\mathop{\Bigm{|}\Bigm{|}}_{h=1}^{H}\sum_{j=1}^{I}\alpha_{i,j}^{h}\cdot\textbf{V}^{h}\textbf{e}_{j};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,j}^{h}=\delta\left(\frac{(\textbf{Q}^{h}\textbf{e}_{i})^{\top}(\textbf{K}^{h}\textbf{e}_{j})}{\sqrt{d/H}}\right)$ (1)
[…]
III-C Heterogeneous Spatio-Temporal GNNs
Our methods begins with designing a spatio-temporal graph $\mathcal{G}$, which is composed of three matrixes, namely region POI matrix $\mathbf{P}$, human trajectory matrix $\mathbf{M}$ and spatio distance matrix $\mathcal{D}$. Thus the relations of the graph $\mathcal{G}$ are classified into five types, namely relations $\mathcal{R}_{p}$ based on POI similarities of regions, relations $\mathcal{R}_{m}$ based on urban flow between regions, relations $\mathcal{R}_{d}$ based on distances of regions, relations $\mathcal{R}_{pm}$ between the POI subgraph and the mobility subgraph, and relations $\mathcal{R}_{md}$ between the mobility subgraph and spatio distance graph. And the relation set $\mathcal{R}$ is shown as $\mathcal{R}=(\mathcal{R}_{p},\mathcal{R}_{m},\mathcal{R}_{d},\mathcal{R}_{pm},\mathcal{R}_{md})$.Thus, relation-aware message passing paradigm is represented as:
$\textbf{h}_{i}^{(l+1)}=\sigma(\sum_{\gamma\in\mathcal{R}}\sum_{j\in\mathcal{N}^{\gamma}_{i}}\alpha_{i,\gamma}\textbf{W}_{\gamma}^{(l)}\textbf{h}_{j}^{(l)});\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,\gamma}=\frac{1}{|\mathcal{N}^{\gamma}_{i}|}$ (2)
where $\mathcal{N}^{\gamma}_{i}$ denotes neighbour set of region $r_{i}$ via relation $\gamma$ ($\gamma\in\Gamma$). $\mathcal{R}$ is the relation set. $\textbf{h}_{i}^{(l+1)},\textbf{h}_{j}^{(l)}\in\mathbb{R}^{d}$ represents the hidden embedding vectors of the $i$-th region vertex in the $(l+1)$-th graph layer and the $j$-th region vertex in the $l$-th graph neural layer, respectively. $\sigma(\cdot)$ denotes the ReLU activation function. $\alpha_{i,\gamma}\in\mathbb{R}$ represents the normalization weight of region vertex pair $(r_{i},r_{j})$, which is calculated using the degrees of the vertices. $\textbf{W}_{\gamma}^{(l)}\in\mathbb{R}^{d\times d}$ presents the learning weights for the $(l)-th$ iteration. STGMAE combines the multi-order embeddings to fully utilize the data obtained from the multi-hop neighborhood. The cross-view message passing mechanism is shown in the following formula:
$\textbf{H}=\sum_{l=0}^{L}\textbf{H}^{(l)};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textbf{H}^{(l+1)}=\sigma(\textbf{D}^{-\frac{1}{2}}\textbf{A}\textbf{D}^{-\frac{1}{2}}\textbf{H}^{(l)}\textbf{W}_{\gamma}^{(l+1)\top})$ (3)
where $\textbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the embedding matrix, whose rows are regional embedding vectors $\textbf{h}_{i}$. $L$ is the number of graph iterations. $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes the adjacent matrix with self-loop. The diagonal degree matrix is denoted as $\textbf{D}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$.
III-D The Layout of Spatial-Temporal Graph Mask Autoencoder (STGMAE)
Given the spatio-temporal graph $\mathcal{G}$, the encoder function $f_{E}$ and the decoder function $f_{D}$, Graph mask autoencoder (GMAE) aims to reconstruct the spatio-temporal graph $\mathcal{G}^{\prime}$ as follows:
$\textbf{H}=f_{E}(\tilde{\mathbf{A}},\tilde{\mathbf{X}}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{G}^{\prime}=f_{D}(\tilde{\mathbf{A}}^{\prime},\tilde{\mathbf{H}})$ (4)
where $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{X}}$ denote the masked adjacent matrix and mask node embedding matrix respectively. And $\tilde{\mathbf{A}^{\prime}}$ and $\tilde{\mathbf{H}}$ represent the remasked adjacent matrix and the remasked node embedding matrix respectively. Details of GMAE design for the spatio-temporal graph $\mathcal{G}$ are in the following sections.
Input: The heterogeneous spatial-temporal graph $\mathcal{G}$, the maximum epoch number $E$, the learning rate $\eta$;
Output: Regional Embedding $\mathbf{H}$
and trained parameters in $\mathbf{\Theta}$;
1 Initialize all parameters in $\mathbf{\Theta}$;
2 Train the framework STGMAE by Equation 9
3 for $epoch=1,2,...,E$ do
4       Random select some nodes and mask them according to Equation 6;
5       Random select some links and mask them;
6       Send the masked nodes features and the adjacent matrix to GCN Encoder;
7       Remark the node features and the adjacent matrix and send them to GCN Decoder;
8       Compute the cosine similarity loss $\mathcal{L}_{1}$ following Equation 8 and MSE loss $\mathcal{L}_{2}$ according to Equation 8;
9       Minimize the loss $\mathcal{L}_{rec}$ by Equation 9 using gradient decent with learning rate $\eta$;
10       for $\theta\in\mathbf{\Theta}$ do
11             $\theta=\theta-\eta\cdot\frac{\partial\mathcal{L}_{rec}}{\partial\theta}$
12       end for
13
14 end for
Return $\mathbf{H}$ and all parameters $\mathbf{\Theta}$;
Algorithm 1 The STGMAE Learning Algorithm
III-D1 Motivation of Features and Structure Reconstruction
The primary motivation behind reconstructing features and structure in GAE lies in the versatility they offer to address different tasks and applications. In certain scenarios, an effective representation of the graph’s structure $\mathbf{A}$ and node features $\mathbf{X}$ can then be utilized for tasks like link prediction or node classification. This can be formulated with coefficient $\alpha$ and activation $\sigma$ as follows:
$\begin{matrix}\underbrace{\mathcal{L}}\\ \text{Graph}\end{matrix}\quad=\quad\begin{matrix}\underbrace{\|\mathbf{X}-\hat{\mathbf{X}}\|_{F}^{2}}\\ \text{Node}\end{matrix}+\quad\alpha\cdot\begin{matrix}\underbrace{\|\mathbf{A}-\sigma(\mathbf{H}\cdot\mathbf{H^{T}})}\|_{F}^{2}\\ \text{Structure}\end{matrix}$ (5)
However, GAE often prioritizes reconstructing the features $\mathbf{X}$ while minimizing or ignoring the reconstruction error related to the structure $\mathbf{A}$. Actually, by adapting $\alpha$ based on the specific task and application, GAEs can effectively learn representations that cater to the demands of diverse graph-based problems. Inspired by recent studies [20, 31] and the significance of different types in our heterogeneous graph, we propose to reconstruct both the structure and features in our STGMAE. Besides, we use a mask mechanism-based generative self-supervised learning paradigm, which allows for effective learning of the spatio-temporal dependencies of region-wise within the graph.
Jointly reconstructing both the structure and features offers several key advantages: (1) Enhanced graph representation. The model can achieve superior graph representation learning and node classification, which is particularly crucial for graphs with multiple relations. (2) Improved relationship nuance capture. It can effectively captures the subtleties of the relationships between nodes, resulting in a more accurate representation of the graph structure. Such a detailed representation is challenging to achieve due to the complexity and multifaceted nature of node relationships.
III-D2 Reconstructing Features and Structures using Mask Mechanism
Developing effective feature-oriented Graph Autoencoders (GAEs) is challenging due to the limited node feature dimensionality in graphs. Existing GAEs, primarily focused on feature reconstruction, often neglect the risk of learning the ”identity function” when the hidden vector dimension surpasses the input, rendering the learned representation useless. To address this issue, we propose the utilization of denoising autoencoders, which intentionally introduce corruption to the input data. Masked autoencoders have demonstrated success in computer vision (CV) and natural language processing (NLP) tasks, where masking is employed as a corruption mechanism. In our suggested approach for GAEs, we advocate for the incorporation of masked autoencoders. Specifically, we randomly select a subset of nodes denoted as $\widetilde{\mathcal{V}}\in\mathcal{V}$ and mask their node features by utilizing a learned vector $x_{M}\in\mathbb{R}^{d}$. This strategy aims to prevent GAEs from learning the ‘identity function’ and enhances the quality of the learned embeddings. By integrating masked autoencoders, we effectively address the challenge of feature reconstruction in GAEs and pave the way for more efficient and meaningful learning of feature-oriented graph representations. The masking process can be defined using Eq. 6:
$x_{i}=\begin{cases}x_{M},&\text{if }v_{i}\in\widetilde{\mathcal{V}}\\ x_{i},&\text{if }v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$ (6)
Given the masked node feature matrix $\widetilde{\mathbf{X}}$, one of the objectives of STGMAE is to reconstruct masked node features of regions (e.g., $r_{i}$) in $\widetilde{\mathcal{V}}$ and masked the adjacent matrix $\widetilde{\mathbf{A}}$ via a learned vector $a_{M}$. In GNNs, the hidden embedding vector of each node is affected by its neighboring nodes’ embedding vectors according to the message passing mechanism. However, selecting nodes via scores can lead to bias, for example, if all neighbors of one node are selected or masked. To prevent bias in node selection, random node selection or sampling is taken into account the entire graph structure rather than just local node properties, which also improves robustness.
III-D3 Remask Decoder via GNNs
The role of the decoder function $f_{D}$ is to map the hidden embedding vectors $\tilde{\mathbf{H}}$ to the input graph $\mathcal{G}^{\prime}$ and its node embedding matrix $\mathbf{X}^{\prime}$, based on the semantic level of the target input graph $\mathcal{G}$ and its corresponding node embedding matrix $\mathbf{X}$ [32]. For example, in NLP, an MLP is typically sufficient for targets of one-hot missing words with rich semantics, as there is no structure semantics to consider. However, in CV, a transformer is often used as the decoder to recover pixel spots with few semantics.
To better capture graph structure information, we adopt GCN as the decoder in our method. This is in contrast to traditional GAEs that typically utilize an MLP to decode the hidden embedding matrix, which results in a similar output to the input graph embedding matrix. Using an MLP as the decoder cannot capture the structure semantics of graphs, so we choose GCN to better capture this.
To further improve the ability of the encoder to learn compressed graph representations, we adopt a remask mechanism in the decoder, inspired by [21]. The process is represented as following:
$\widetilde{h}_{i}=\begin{cases}h_{M}&\leavevmode\nobreak\ \leavevmode\nobreak\ v_{i}\in\widetilde{\mathcal{V}}\\ h_{i}&v_{i}\notin\widetilde{\mathcal{V}}\end{cases}$ (7)
Here, $\widetilde{\mathbf{H}}=REMASK(\mathbf{H})$, where a masked node is forced to build its input feature based on the neighboring non-masked nodes’ hidden embedding vectors, under the guidance of the GCN decoder. This approach helps ensure that the learned embeddings are both informative and compressed, which can improve the overall region representation ability of our framework STGMAE.
III-E Model Optimization and Complexity
To optimize our method STGMAE, we adopt the cosine similarity loss, denoted as $\mathcal{L}_{1}$, to optimize remasked node features and initial input node features. Besides, to help optimize structure masking, another loss MSE is adopted. These two losses are shown as:
$\mathcal{L}_{1}$ $=\frac{1}{N}\sum_{i=1}^{N}\left(1-\frac{h_{i}^{T}\,\widetilde{h}_{i}}{\|\widetilde{h}_{i}\|\cdot\|h_{i}\|}\right);$ (8)
$\mathcal{L}_{2}$ $=\left\|\mathbf{E}-\sigma\left(\widetilde{\mathbf{A}}^{\prime}\cdot\widetilde{\mathbf{A}}^{\prime T}\right)\right\|_{2}^{2}$
where $\mathbf{E}$ is the hidden matrix of initial adjacent matrix $\mathbf{A}$.
By leveraging generative self-supervised learning tasks during pretraining, our proposed framework, STGMAE, aims to produce high-quality region embeddings. The optimization goal of our framework STGMAE can be formalized as follows:
$\min\limits_{\theta}\mathcal{L}_{rec}(\mathbf{A},\mathbf{X},\widetilde{\mathbf{A}}^{\prime},\widetilde{\mathbf{X}}^{\prime}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{L}_{rec}=\mathcal{L}_{1}+\mathcal{L}_{2}$ (9)
where $\theta$ denotes the parameters of the model, $\mathcal{L}_{rec}$ is the reconstruction loss that measures the difference between the original and reconstructed adjacency matrix $\mathbf{A}$ and node feature matrix $\mathbf{X}$. The objective function ensures that the model can accurately reconstruct both the graph structure and node features, while the second term encourages the model to learn representations that are useful for predicting node features in a self-supervised manner. Overall, the optimization goal of our framework reflects our aim to learn high-quality region embeddings that can be used for downstream tasks such as node classification, link prediction, and community detection. And the time complexity of our framework STGMAE is divided into two parts: first of all, the multi-view space-time message passing paradigm has a complexity of $\mathcal{O}(|\mathcal{V}|\times L\times d)$, where $N$ and $L$ denote the number of edges and graph propagation layers, respectively. Secondly, for loss calculation, STGMAE takes $\mathcal{O}(S\times|\tilde{\mathcal{V}}|\times d)$, where $S$ denotes the number of sampled nodes included in each batch for the spatio-temporal graph.
III-F In-depth Theory Analysis of STGMAE
In this section, we aim to provide in-depth analysis of our framework STGMAE. We present a small STGMAE loss denoting a small spatio-temporal alignment loss. Firstly, we have the following assumption: There exists a pseudo-inverse encoder $f_{E}$ which also the resulting pseudo autoencoder $s_{g}=\left\{g,f_{g}\right\}$satisfies $\mathbf{E}||s_{g}(z)-z||\leq\epsilon$, where $z$ denotes masked representation of the input graph $\mathcal{G}$. Following this assumption, our method performs input-output alignment on the masked spatio-temporal graph. The lower bound of $\mathcal{L}_{rec}\geq\mathcal{L}_{asym}-\epsilon+C$, where $C$ is a constant number and $\mathcal{L}_{asym}$ is presented as following for the spatio-temporal graph $\mathcal{G}$:
$\mathcal{L}_{asym}=-\mathbf{E}_{x_{1},x_{2}}f(x_{1})s_{g}(x_{2})=-tr(\mathbf{H}_{g}^{T}\bar{\mathbf{A}}_{m}\mathbf{H})$ (10)
where $\mathbf{H}$ is the output matrix of our method STGMAE on $\mathcal{G}_{1}$ whose $x_{1}$-row is $\mathbf{H}_{x_{1}}=\sqrt{d_{x_{1}}}f(x_{1})$, and $\mathbf{H}_{g}$ is the output matrix of $s_{g}$ on $\mathcal{G}_{2}$ whose $x_{2}$-th row is $(\mathbf{H}_{g})_{x_{2}}=\sqrt{d_{x_{2}}}s_{g}(x_{2})$. Thus, a small STGMAE loss implies a small alignment loss as a lower bound, which also verifies align mask and unmask view via STGMAE. This again verifies that our method can obtain the same level performance as the contrastive learning loss without data augmentation.
Table I: Data Description of Experimented Datasets
Data Description of Chicago Data Description of NYC data
Census
Boundaries of 234 regions
split by streets in
a certain district, Chicago
Boundaries of 180 regions
split by streets in
Manhattan, New York
Taxi trips
Total 386,272 taxi
trips during a month
Total 1,445,285 taxi
trips during a month
Crime data
Total 321,876
crime records during 1 year
Total 108,575 crime
records during 1 year
POI data
Total 3,680,125 POI locations
of 130 categories
Total 20,569 POI locations
of 50 categories
House price
Total 44,447 house price data
in a certain district, Chicago
Total 22,540 house price data
in Manhattan, New York
IV Evaluation
We conduct experiments to answer the following questions: RQ1: How does the STGMAE fare in terms of different baselines across multiple spatial-temporal learning applications including traffic prediction, crime prediction and house price prediction? RQ2: How do various data perspectives and generative elements impact the effectiveness of region representation? RQ3: How well does our STGMAE fare in representation learning over areas with various levels of data sparsity? RQ4: What effects do different hyperparameter settings have on the performance of the STGMAE’s region representation? RQ5: How the efficiency our method STGMAE than other region representation methods (e.g., MV-PN, MVURE and MGFN)? RQ6: What advantages do our STGMAE using learned representations across regions have?
IV-A Experimental Setup
IV-A1 Datasets and Protocols
With the aim of assessing the effectiveness of our STGMAE framework, we gathered real-world datasets from two major cities, namely Chicago and New York City. These datasets were utilized to perform three spatio-temporal mining tasks: crime prediction and traffic flow forecasting, as well as property price prediction. To ensure a comprehensive evaluation, we included various types of crimes in our datasets. For the Chicago database, crimes such as Theft, Battery, Assault, and Damage were considered, while the NYC database encompassed crimes such as Burglary, Larceny, Robbery, and Assault. The selection of crime types aligns with the parameters in [33]. A detailed description of the data is shown in Table I.
IV-A2 Hyperparameter Settings
Following existing region representation studies [27, 24], we set the dimensionality $d$ as 96. And from the hyperparameter study experiments, we find that the framework STGMAE obtains the best performance when the GCN depth is set as 2. And the learning rate is 0.001 with weight decay 0.0005. For crime prediction backbone model, the learning rate of ST-SHN is set as 0.001 and weight decay is 0.96. The depth of the spatial aggregation layer is set as 2. For traffic prediction backbone model ST-GCN, the input length is 12 time intervals with each interval setting as 5 minutes. And the output is 15 minutes’ results of traffic flow prediction. Most of baselines are utilized following their released codes.
IV-A3 Implementation Details
To be fair, we implement STGMAE and other baselines in Python 3.8, pytorch 1.7.0 with cuda 11.3 for our method implementation and tensorflow 1.15.3 for traffic prediction and crime prediction. The experiments are conducted on a server with 10-cores of Intel(R) Core(TM) i9-9820X CPU @ 3.30GHz 64.0GB RAM and one Nvidia GeForce RTX 3090 GPU.
Refer to caption Refer to caption
Figure 2: Ablation study of STGMAE on traffic prediction
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Ablation study of STGMAE on crime prediction
Table II: Overall performance in crime forecasting, traffic prediction, and house price prediction.
Model Crime Prediction Traffic Prediction House Price Prediction
CHI-Crime NYC-Crime CHI-Taxi NYC-Bike NYC-Taxi CHI-House NYC-House
MAE $\downarrow$ MAPE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$ MAE $\downarrow$ MAPE $\downarrow$
ST-SHN 2.0259 0.9987 4.4004 0.9861
ST-GCN 0.1395 0.5933 0.9240 1.8562 1.4093 4.1766
Node2vec 1.6334 0.8605 4.3646 0.9454 0.1206 0.5803 0.9093 1.8513 1.3508 4.0105 13137.2178 44.4278 4832.6905 19.8942
GCN 1.6061 0.8546 4.3257 0.9234 0.1174 0.5707 0.9144 1.8321 1.3819 4.0200 13074.2121 42.6572 4840.7394 18.3315
GAT 1.5742 0.8830 4.3455 0.9267 0.1105 0.5712 0.9110 1.8466 1.3746 4.0153 13024.7843 43.3221 4799.8482 18.3433
GraphSage 1.5960 0.8713 4.3080 0.9255 0.1196 0.5796 0.9102 1.8473 1.3966 4.0801 13145.5623 44.3167 4875.6026 18.4570
GAE 1.5711 0.8801 4.3749 0.9343 0.1103 0.5701 0.9132 1.8412 1.3719 4.0337 13278.3256 42.3221 4896.9564 18.3114
GraphCL 1.2332 0.6293 3.3075 0.6771 0.0812 0.5364 0.8582 1.8180 1.3022 3.7029 10752.5693 28.8374 4562.7279 11.3055
RGCL 1.1946 0.6081 3.1026 0.6323 0.0797 0.5052 0.8319 1.8107 1.2973 3.6865 10673.3289 27.5279 4439.0733 10.2375
POI 1.3047 0.8142 4.0069 0.8658 0.0933 0.5578 0.8892 1.8277 1.3316 3.9872 12045.3212 33.5049 4703.3755 16.7920
HDGE 1.3586 0.8273 4.2021 0.7821 0.0865 0.5502 0.8667 1.8251 1.2997 3.9846 11976.3215 30.8451 4677.6905 12.5192
ZE-Mob 1.3954 0.8249 4.3560 0.8012 0.1002 0.5668 0.8900 1.8359 1.3314 4.0366 12351.1321 38.6171 4730.6927 16.2586
MV-PN 1.3370 0.8132 4.2342 0.7791 0.0903 0.5502 0.8886 1.8313 1.3306 3.9530 12565.0607 39.7812 4798.2951 17.0418
CGAL 1.3386 0.7950 4.1782 0.7506 0.1013 0.5682 0.9097 1.8557 1.3353 4.0671 12094.5869 36.9078 4731.8159 16.5454
MVURE 1.2586 0.7087 3.7683 0.7318 0.0874 0.5405 0.8699 1.8157 1.3007 3.6715 11095.5323 34.8954 4675.1626 15.9860
AutoST 1.0480 0.4787 2.1073 0.5241 0.0665 0.4931 0.8364 1.8145 1.2871 3.6446 10463.7715 25.7575 4517.7276 7.1660
MGFN 1.2538 0.6937 3.5971 0.7065 0.0831 0.5385 0.8783 1.8163 1.3266 3.7514 10792.7834 29.9832 4651.3451 12.9752
STGMAE 1.0438 0.4608 2.1060 0.5203 0.0606 0.4817 0.7787 1.8079 1.2778 3.6335 10375.9982 23.1751 4432.1934 6.8785
P-Value 5.27e-12 1.74e-14 2.7e-13 8.22e-11 1.32e-6 7.34e-10 2.13e-10 9.38e-4 1.24e-18 5.37e-21 2.18e-5 1.78e-4 1.91e-4 2.75e-4
Table III: Overall performance comparison in crime prediction on both Chicago and NYC datasets.
Chicago New York City
Model Theft Battery Assault Damage Burglary Larceny Robbery Assault
MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE
Node2vec 1.1378 0.9862 1.7655 0.8970 1.9631 0.9714 1.9015 0.9657 4.9447 0.8092 0.7272 0.6532 1.0566 0.8040 1.2411 0.9967
GCN 1.1065 0.9643 1.3012 0.8094 1.5431 0.8094 1.5031 0.8056 4.6993 0.7912 0.49994 0.4178 1.0655 0.8004 1.2407 0.9890
GAT 1.1123 0.9759 1.3215 0.8344 1.5892 0.8241 1.5387 0.8277 4.7055 0.7944 0.5023 0.4019 1.0653 0.8027 1.2403 0.9949
GraphSage 1.1231 0.9790 1.3574 0.8561 1.6016 0.8563 1.5761 0.8432 4.7313 0.8066 0.5213 0.4314 1.0719 0.8110 1.2418 0.9965
GAE 1.1043 0.9614 1.3065 0.7984 1.5379 0.7914 1.4986 0.8033 4.7013 0.7910 0.5012 0.4289 1.0679 0.8012 1.2405 0.9958
GraphCL 0.9994 0.8992 1.0578 0.6174 1.2938 0.5683 1.2457 0.6181 4.5047 0.7288 0.3834 0.3091 1.0498 0.5961 0.7724 0.6929
RGCL 0.9377 0.8779 1.0356 0.5865 1.2764 0.5457 1.2336 0.5975 4.4712 0.7064 0.3652 0.2832 0.9987 0.5769 0.7512 0.6743
POI 0.9733 0.9341 1.1065 0.7513 1.4089 0.7541 1.4076 0.7697 4.6939 0.7825 0.4969 0.4172 1.0660 0.7970 1.2400 0.9943
HDGE 0.9545 0.9012 1.0887 0.7389 1.3970 0.7217 1.3768 0.7349 4.5658 0.7160 0.4734 0.3930 1.0507 0.6731 1.1551 0.9870
ZE-Mob 1.0983 0.9547 1.3142 0.8236 1.5345 0.8163 1.5138 0.8273 4.7570 0.8013 0.5186 0.4307 1.0722 0.8093 1.1403 0.9975
MV-PN 0.9613 0.9146 1.0946 0.7452 1.4013 0.7393 1.3582 0.7218 4.6329 0.7502 0.4213 0.3708 1.0642 0.7840 1.1091 0.9982
CGAL 0.9589 0.9014 1.0897 0.7403 1.3995 0.7345 1.3698 0.7296 4.6013 0.7203 0.4113 0.3651 1.0714 0.7765 1.1009 0.9894
MVURE 0.9365 0.8910 1.0631 0.6957 1.3709 0.6375 1.3037 0.6567 4.5907 0.7144 0.4077 0.3262 1.0578 0.5889 0.8410 0.6943
MGFN 0.9231 0.9015 1.0804 0.5824 1.3016 0.6072 1.2563 0.6503 4.5646 0.7994 0.4285 0.3084 1.0475 0.6310 0.8319 0.7096
STGMAE 0.7860 0.7365 1.0377 0.4207 1.1830 0.4214 0.9252 0.4596 4.3095 0.6404 0.2958 0.1924 0.8788 0.3789 0.6654 0.5537
IV-A4 Baselines
We use 15 baselines from three research lines namely, graph representation approaches, graph contrastive learning models, and spatial-temporal region representation methods for comparison in order to thoroughly assess our STGMAE model.
Graph Representation Approaches:
  • Node2vec [34] encodes graph structural data using a random walk-based skip-gram.
  • GCN [28] carries out the convolution-based message transfer along the edges between neighboring nodes for refinement.
  • GraphSage [35] is a graph neural architecture that permits the accumulation of information from the sub-graph structures.
  • GAE [17] maps node into latent embedding space by the Graph Auto-encoder using the input reconstruction objective.
  • GAT [36] varies the degrees of significance among nearby nodes, and improves GNNs’ capacity to discriminate between diverse inputs. To optimize the performance of GNN-based approaches, the number of information propagation layers is adjusted from a range of {1,2,3,4}.
Graph Contrastive Learning Models:
  • GraphCL [37] is based on the maximizing of mutual knowledge and generates many contrastive viewpoints for augmentation. The goal is to ensure embedding consistency across various connected views via the InfoNCE loss.
  • RGCL [38] does data augmentation based on the created rationale generator, is used to do so. Different GCL models’ temperature parameters have been fine-tuned to manage the contrastive augmentation in order to allow for fair comparison.
Spatial-Temporal Region Representation Methods:
  • POI [29] uses TF-IDF tokenization with the provided POI matrix and the POI properties to represent spatial regions.
  • HDGE [22] creates a crowd flow graph using human trajectory data and embeds regions into latent vectors to protect the structural data of the network and capture mobility patterns.
  • ZE-Mob [23] uses region correlations to create embeddings while taking into account human movement and taxi moving traces based on multi-view data sources.
  • MV-PN [26] models intra-regional and inter-regional correlations with an encoder-decoder network.
  • CGAL [25] is a technique for graph-regularized adversarial learning that uses pairwise graph relationships to embed regions in latent space via the graph attention neural network.
  • MVURE [24] makes use of the graph attention mechanism to simulate region correlations with the built-in properties of the region and data on human mobility.
  • AutoST [15] is a method that utilizes contrastive self-supervised learning method via multi-view graph to model region representation learning problem on traffic and crime prediction tasks.
  • MGFN [27] aggregates information for both intra-pattern and inter-pattern patterns, and encodes region embeddings with multi-level cross-attention without region function information.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Results on NYC and CHI crime for four crime types w.r.t different data density degrees.
IV-B Effectiveness Evaluation (RQ1)
Table II shows a performance comparison of all methods on three tasks (crime prediction, traffic prediction, and house price prediction) in terms of MAE, MAPE, and RMSE.
Crime prediction. We show crime prediction results in Table II and type-specific results in Table III. From Table II, we summarize observations as follows:
  • On both datasets, our framework STGMAE performs best at predicting crimes, indicating the viability of our generative spatio-temporal graph self-supervised learning approach. In specific, we attribute the benefits to: i) STGMAE can capture both the mutli-view area relationships from multi-typed data sources (i.e., POI semantics, human motion traces, geographical coordinates) thanks to masking different types of relations. ii) The adaptive data augmentation way of STGMAE via the mask mechanism on nodes and relations improves the representation ability against noise disturbance and skewed data distribution.
  • While existing approaches for region representation (such as CGAL, MVURE, and MGFN) aim to capture the mobility-based region correlations, they are susceptible to noisy edges in their built mobility-based region graphs. The performance of region representation learning will be hampered by the propagation of information propagation between less-relevant regions. Besides, the spatial dependence modeling in such baselines is constrained by the skewed distribution of mobility data. Our STGMAE   which is complementary to the mobility-based region connections, adds region self-discrimination supervision signals to the region embedding paradigm to solve these limitations.
  • The large enhancement of STGMAE over the compared graph embedding techniques (i.e., GCN and GAT) further supports the significance of generative self-supervised learning paradigm. When a message passes over the edges of a task-irrelevant region in the graph neural paradigm, noise effects will be enhanced. The information compiled from all the nearby mobility-based or POI-based graph regions may be misleading when it comes to encoding the actual underlying crime trends.
Traffic prediction. We present traffic prediction results in Table II with the following observations:
  • The significance of the generative self-supervised learning with the mask mechanism is further supported by the significant performance improvement of STGMAE over the compared graph embedding techniques (such as GCN, GAT, and GAE). In the graph neural paradigm, noise effects will be amplified when messages crosses edges of regions that are unrelated to the task.
  • To find traffic patterns or mine human behavior, region representation methods (i.e., MVURE, MGFN) need enough data, which means they cannot handle long-tail data. Contrarily, contrastive learning with masking nodes or links mechanism has better skewed data distribution adaption.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Hyperparameter study on traffic prediction
IV-C Ablation Study (RQ2)
To evaluate the effectiveness of each component of our framework STGMAE, we perform an ablation study and present results w.r.t traffic prediction and crime prediction in Figure 2 and Figure 3. Particularly, we conduct an ablation study on “RP GCN” by replacing GCN with MLP, “w/o Structure Mask”, “w/o Node Mask”, and “w/o Edge Type Mask” by removing masking links or masking nodes.
Analysis on Crime prediction. The figure demonstrates the impact of different components on region representation for crime prediction. “RP GCN” replacing GCN with a 2-layer MLP, contributes the most to performance improvement compared to “w/o Structure Mask”, “w/o Edge Type Mask”, and “w/o Node Mask” GCN captures spatial-temporal graph structure semantics, unlike MLP. “w/o Node Mask” has a larger impact than “w/o Structure Mask” as masking nodes directly affects region representations. Masking different link types positively affects region representations, emphasizing the significance of masking links in the heterogeneous graph. “w/o Edge Type Mask” which doesn’t mask edges, shows a notable performance decline, highlighting the effectiveness of edge type masking in our STGMAE model. Overall, integrating GCN, node masking, and edge-type masking enhances the performance of our STGMAE framework.
Analysis on Traffic prediction. Based on Figure 2, GCN has the greatest impact on improving region representations and traffic prediction, especially for longer time horizons. Masking edge types significantly improves traffic flow prediction accuracy. Masking nodes and graph structure also contribute positively by refining region representations and capturing spatial relationships. Incorporating POI information enhances region representations and enables the encoding of different region relations. Overall, the combination of GCN, edge type masking, node masking, and POI integration contributes to the success of our STGMAE method in learning region representations and achieving accurate traffic predictions.
Type-Specific Crime Prediction Results. We present a comprehensive evaluation of various criminal offense types in the cities of Chicago and New York. The evaluation measures the performance of 16 different approaches in terms of MAE and MAPE. The foundational method for the other 16 approaches is ST-SHN. The results, shown in Table III, clearly demonstrate that our STGMAE approach consistently outperforms all other methods across all crime categories in both cities. This highlights the significant advantages that our region’s embedding learning framework brings to the table. We attribute the success of our approach to the effectiveness of the multi-view region graph’s graph encoding, which extracts valuable regional features for region representation. Additionally, the generative contrastive method, utilizing masking nodes and different types of links in the heterogeneous graph, plays a crucial role. This further validates the effectiveness of generative self-supervised learning in spatio-temporal graph-related applications such as traffic flow prediction, crime prediction and house prediction, leveraging the mask mechanism on nodes and structures.
IV-D More Effectiveness Evaluation Results (RQ1)
Results for the data on home prices in Chicago and New York City are shown in Table II and more analysis details are below. A Lasso regression approach uses the input embeddings from the learned region representations created using various methods.
  • STGMAE consistently performs at the highest level. On the other hand, MVURE and MGFN perform better when using data from the housing markets of New York and Chicago, respectively. This is because they mine human patterns, which connect comparable or nearby places. We also notice that the POI technique performs well because it creates connections between POIs of regions and core shopping regions, which include POIs that are distinct from those in remote towns.
  • Methods taking advantage of region representations outperform those using network embeddings in terms of performance. The basis for this investigation is that region-wise dependencies from both the spatial and temporal dimensions cannot be effectively encoded by conventional graph-based approaches (such as GCN).
  • Table IV shows results of improvement of existing widely used traffic prediction methods (e.g., GMAN, STGCN) based on several region representation baselines and our method STGMAE. The superior performance further validates the effectiveness of generative self-supervised learning with a masking mechanism for denoising data. In contrast, methods like MVURE and MGFN require significantly more data to uncover human patterns.
Table IV: Overall Performance Comparison in Traffic Prediction Improvement
Methods NYC-Taxi Methods NYC-Taxi
MAE RMSE MAE RMSE
GMAN 2.0812 3.0215 STGCN 1.4093 4.1766
w/MVURE 1.8795 2.8337 w/MVURE 1.3007 3.6715
w/MGFN 1.9261 2.8337 w/MGFN 1.3266 3.7514
w/STGMAE 1.8403 2.7032 w/STGMAE 1.2415 3.6375
IV-E Model Robustness Study (RQ3)
We also conduct experiments to test how well our framework STGMAE stands up to data sparsity. To accomplish this, we assess the forecast accuracy of regions with various densities independently. Here, the ratio of non-zero elements (crime occurs) in the region-specific crime occurrence sequence $\mathbf{Z}_{r}$ is used to estimate the density degree of each region. We divide sparse areas with a crime density degree of less than 0.5 into two groups, (0.0, 0.25] and (0.25, 0.5] specifically. Figure 4 presents the evaluation findings. We see that our STGMAE constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (such MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (such as GNN-based approaches including GCN, GAE and GAT) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
The sencond line of Figure 4 present more details about the evaluation findings of the model robustness. We see that STGMAE constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (i.e., GNN-based approaches including GCN, GAE and GAT) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
IV-F Hyperparameter Study (RQ4)
We conduct the hyperparameter study on the parameters including the mask ratio of nodes and links with the range $[0.1,0.3,0.5,0.7]$, and the number of CGN layer from the range $[2,3,4,5]$ on crime prediction task and traffic predition  w.r.t New York and Chicago in terms of MAE, MAPE and RMSE. We provide results in Figure 5. From figures, we can have the following observations: (1) For the impact of mask ratio of nodes and structure, we find that the model STGMAE can achieve the best performance on crime prediction and traffic prediction on all datasets when the mask ratio of nodes and links is set as 0.7. And the performance decreases with the decrease of mask ratio. This verifies that masking nodes and links do not always bring better model representation ability. Smaller mask ratios (e.g., $[0.1,0.3,0.5]$) may not denoise nosiy nodes and links completely, which leads to bad performance. (2) For the effect of the number of GCN layer, we find that our framework STGMAE achieves the best performance when the number of GCN layer is set as 2. We notice that the performance decreases with the increasing number of GCN layer. This may be caused by over-smoothing issue on region representation with the increasing number of the GCN layer. (3) We also investigate the consuming time of our model STGMAE with varied hyperparameters. We observe that time cost increases with increasing mask ratio of nodes and links, since our model STGMAE requires to reconstruct more nodes’ representation and new links when masking more nodes and links. Similarity, time consuming is larger when the more layers of GCN due to that the message passing between GCN layers need more time.
IV-G Model Efficiency Study (RQ5)
We investigate the efficiency of our method STGMAE and existing region representation learning methods  w.r.t training time. We provide results in Table V. All of these methods are adopted in python 3.8, pytorch 1.7.0 with cuda 11.3 and tensorflow 1.15.3 (GPU version). The experiments are conducted in the same server. From the table, By comparing the model efficiency of STGMAE to the baselines (e.g., HDGE, MVURE and MGFN), we can see that it is competitive, indicating that it has the capacity to handle massive amounts of spatial-temporal data. The performance of the region representation is greatly improved by the generative self-supervised learning paradigms while incurring little additional processing expense.
Table V: Computational cost (seconds) of our STGMAE and SOTA spatial-temporal region representation methods. MAE and MAPE are estimated for NYC crime prediction results.
Models HDGE ZE-Mob MV-PN CGAL MVURE MGFN Ours
Training 303.7 82.7 33.4 4144.8 240.7 852.3 124.5
MAE 4.2021 4.3560 4.2342 4.1782 3.7683 3.5971 2.1060
MAPE 0.7821 0.8012 0.7791 0.7506 0.7318 0.7065 0.5203
Figure 6: Case study of STGMAE on New York City. Our STGMAE can alleviate the connection bias in spatial-temporal graphs by refining the salient relational signals between far away regions (region 1 & region 124).
IV-H Case Study (RQ6)
In this part, we use a case study to demonstrate how well our STGMAE can understand the relationship between global regions and geographical semantics in Figure 6. We sample 1 region pairings, or distant regions, specifically (region-1, region-124). Even though regions 1 and 158 are geographically far apart, as seen in the provided Figure, their latent meanings are well preserved in our encoded embeddings. Their cluster results from the MGFN baseline, however, demonstrate a significant difference. In contrast, the region clusters discovered using our STGMAE method exhibit similarity. These findings show that our STGMAE has the advantage of incorporating global region dependencies over the entire metropolitan landscape, which is important in real-life applications.
V Conclusion
We introduce a novel generative spatial-temporal graph self-supervised learning paradigm for region representation. Our approach incorporates mask mechanisms to address limitations in existing area representation methods. Through extensive experiments on real-world datasets and three spatial-temporal mining tasks, we demonstrate the effectiveness of our STGMAE in learning area graph representations from multi-view spatial-temporal data. Our contributions include proposing a self-supervised learning paradigm using generative models and masks, addressing challenges in area representation methods, and achieving superior performance compared to state-of-the-art approaches in various spatial-temporal mining tasks, such as crime prediction.
VI Acknowledgement
This project is partially supported by HKU-SCF FinTech Academy and Shenzhen-Hong Kong-Macao Science and Technology Plan Project (Category C Project: SGDX202108 23103537030) and Theme-based Research Scheme T35-710/20-R. We appreciate the help from Dr. Huang and Dr. Xia.